Jimmie cone has one container of 25% butterfat vanilla ice cream and another of 50% butterfat. How much of each should they use to produce 1,000 gallons of vanilla ice cream that is 45% butterfat?
Answer by macston(5194)

.
X=gallons of 25%; Y=gallons of 50%
.
X+Y=1000
Y=1000-X
.
0.25X+0.50Y=0.45(1000)
0.25X+0.50(1000-X)=450
0.25X+500-0.50X=450
-0.25X=-50
X=200 ANSWER 1: They should use 200 gallons of 25%.
.
Y=1000-x=1000-200=800 ANSWER 2: They should use 800 gallons of 50%.
.
CHECK:
.
0.25X+0.50Y=0.45(1000)
0.25(200)+0.50(800)=0.45(1000)
50+400=450
450=450